# Towards the Usage of Window Counting Constraints

### v2

@@ 4. Example and Experimental Results
Figure 8. Two-player game graph $A$ for a safety game. States represented as circles are controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. $\Sigma_{E}=\{a\}$ and outgoing transitions of states controlled by EGOEGO without label mean that $EGO$ plays $\neg a$ in those moves. The played letters of $ALTER$ are not important in this example, hence labels of outgoing transitions of ALTERALTER-controlled states are omitted. To keep the example simple, we consider a safety game without unsafe states, that is, $\text{Win}=\mathtt{Safe(R)}$ for $R$ containing all states of $A$. EGOEGO shall fulfill the counting constraint $C=CC_{min}(EGO,a,1,7)$ (EGOEGO plays $a$ at least one time in 7 turns), making $G=(A,\text{Win},\{C\})$ a game with counting constraints.
@@
  • For increasing the counting constraint length for the next iteration, we implemented the sequential and the alternating mode explained above. Each game is played in each of the available modes.
As comparison, the same base implementation is used, but the first (and only) increment is done directly with the full length of all counting constraints, which translates to synthesis without incrementing and direct encoding of the required memory for the full constraint length into the game graph. Table 1 summarizes the performance results of the presented incremental synthesis approach. Each game was solved three times: first in sequential mode for increasing the constraints, second in alternating mode and the third time without incrementing (by going directly to the full constraint lengths). For each of those runs of the synthesis algorithms, the required computational time and the number of states in the last considered situation graph is presented in the table. Additionally, it is noted how many constraints on $EGO$ were given and if a winning strategy for $EGO$ exists. A question mark “?” in the table stands for data that was not recorded due to hardware limitations (thrashing due to reached memory limits; in fact, this phenomenon could also have lead to the high computational time in experiment number 2 without incrementing).
@@
sequential alternating
No. Win? # $\mathcal{C}$ # states time [sec] # states time [sec] # states time [sec]
1 Yes 1 234,477 28 234,477 31 3,189,957 253
@@ 5 Yes 2 537,970 111 152,882 51 ? ?
The realized comparisons show the great potential of successively enlarging counting constraints, allowing for incomplete graph constructions due to retrieved information on already winnable states of prior graphs instead of encoding the full constraints directly in a graph for strategy synthesis. In all of the experiments, the incremental synthesis required significantly less time and memory. Note that the number of states in Table 1 refers to the number of states in the last constructed situation graph. If incrementing is used, additional memory for the states of the already calculated winning regions of previous increments is required. It suffices to remember a minimal set $M$ of states, such that each state of any of the already calculated winning regions is an extension of a state in the minimal set $M$. The number of states in $M$ can be larger than the next situation graph (as can be seen in the example above and this is also the case for experiment 3 in the table). However, the set $M$ is (heuristically) much smaller than the situation graph without incrementing. For example, experiment 3 in Table 1 required to remember 68,937 states of winning regions for sequential increments and 8,401 states for alternating increments, which is orders of magnitude smaller than the 1,723,345 states in the situation graph without incrementing.
Direct synthesis without increments will outperform synthesis with increments if the winning region is empty for all increments but the one with full length of all counting constraints, since incrementing will then only add a non-beneficial overhead. However, it is usually not known beforehand whether the winning regions of all but the last increment will be empty.
@@ 5. Discussion and Future Work
Towards cooperative games: The presented approach did not consider incrementing over constraints of $ALTER$, but handled them as being fixed. The rationale for this lies in the property ALTERALTER cannot be forced into constraint violations” of games with counting constraints (Definition 2). In general, a game with counting constraints may satisfy the requirement of $ALTER$ always being able to adhere to its counting constraints, only to find the requirement violated for the game with a modified counting constraint as used in the increments. An illustrative example is provided in Figure 10. $ALTER$ has the counting constraint $CC_{min}(ALTER,b,1,3)$, i.e. $ALTER$ plays $b$ at least once in three of its turns. Recall that ALTERALTER cannot be forced into constraint violations” is defined as $ALTER$ is able to enlarge each prefix that satisfies the constraint such that the resulting prefix is also satisfying the constraint. This property is fulfilled when considering the game graph and the constraints $CC_{min}(ALTER,b,1,1)$ or $CC_{min}(ALTER,b,1,3)$. However, it is violated for $CC_{min}(ALTER,b,2,3)$, since the prefix $(1,a,2,\neg b,4,a,5)$ satisfies the constraint [^1] ALTERALTER could play $b$ forever to complete the prefix to an infinite play that satisfies the constraint. This play is not in $G$, but nonetheless is sufficient according to the definition of a prefix satisfying a constraint in Section 2., but there is no possibility for $ALTER$ to still satisfy the constraint with the next turn. Since the definition of a winning strategy relies on the game property of $ALTER$ not being forceable into counting constraint violations, 1 cannot be extended to increments over $ALTER$-constraints. However, such an extension would offer additional potential for more efficient synthesis algorithms.
@@
Symbolic representation: The presented synthesis approach uses an explicit representation of states in the situation graph as arena. However, symbolic synthesis proves to be significantly more efficient than explicit synthesis algorithms for many (but not all) applications [Fin16]. Since the presented approach already has similarities to antichains and the states of the arena have a special structure (representing a snippet of the history of a play), we expect that the approach can be transformed into a symbolic algorithm. We plan to investigate a symbolic version of the algorithm and to compare its performance with its explicit version.
Symmetric synthesis: In many application scenarios, the considered system is a system of systems, whose individual systems are identical in their hardware and run in the same context. From the development, production and maintenance point of view, it is preferable to also have the same software and in particular the same controller for all of those systems, as it is typically realized in industrial applications. An example is a fleet of transport robots that take over logistic tasks on a factory floor. Some synthesis approaches that consider such symmetric synthesis (e.g., [EF17], [CS11], [RCB24]) exist, but to the best of our knowledge, there still remain open gaps. We are especially interested in such symmetric setting because (1) it is of high relevance for the same application domain where the problem of synthesis with counting constraints arose (fleet of transport robots) and
@@ 6. Conclusion
Synthesis algorithms for reactive systems are promising tools for various engineering tasks, most prominently for the creation of correct-by-construction controllers and for checking the feasibility of specifications. The efficiency of such algorithms is a challenge for getting synthesis into application, since the translation of the system specification into an automaton that is suitable for synthesis is costly in terms of memory and computational time. The exploitation of specific properties in the specification can help to overcome this challenge. In this article, we have shown the potential of incremental synthesis algorithms for specifications with monotonicity properties as for the presented counting constraints. With each increment, the automaton encoding the specification is becoming larger. The key idea is to gather information in each increment that can be used in the next increment to reduce the size of the automaton. The precise nature of this information depends on the considered specification. We have shown an incremental algorithm for a set of counting constraints of the form “the system does a specific move $m$ at least/at most in $k$ turns out of $l$, in which information on winnable states of the automaton in one increment can be used to determine which parts of the automaton for the next increment do not need to be constructed. The approach allows to tackle games with various winning conditions and the integration of specialized synthesis algorithms for calculating the winning regions of intermediate games occurring in the incrementing process. In the presented experiments, the incremental approach requires significantly less memory and computational time than direct synthesis with full specification translation into one automaton. As future work, we plan to extend the incremental synthesis in four dimensions: (1) Consideration of more cooperative behavior between system and its environment instead of a purely adversarial setting, (2) identification of new specification types with monotonicity properties that can be exploited via incremental synthesis, (3) transformation of the synthesis approach to a symbolic synthesis algorithm and (4) consideration of systems of systems, in which identical systems follow the same strategy.

### v3

@@ 4. Example and Experimental Results
Figure 8. Two-player game graph $A$ for a safety game. States represented as circles are controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. $\Sigma_{E}=\{a\}$ and outgoing transitions of states controlled by EGOEGO without label mean that $EGO$ plays $\neg a$ in those moves. The played letters of ALTERALTER are not important in this example, hence labels of outgoing transitions of ALTERALTER-controlled states are omitted. To keep the example simple, we consider a safety game without unsafe states, that is, $\text{Win}=\mathtt{Safe(R)}$ for $R$ containing all states of $A$. EGOEGO shall fulfill the counting constraint $C=CC_{min}(EGO,a,1,7)$ (EGOEGO plays $a$ at least one time in 7 turns), making $G=(A,\text{Win},\{C\})$ a game with counting constraints.
@@
  • For increasing the counting constraint length for the next iteration, we implemented the sequential and the round-robin mode explained above. Each game is played in each of the available modes.
As comparison, the same base implementation is used, but the first (and only) increment is done directly with the full length of all counting constraints, which translates to synthesis without incrementing and direct encoding of the required memory for the full constraint length into the game graph. Table 1 summarizes the performance results of the presented incremental synthesis approach. Each game was solved three times: first in sequential mode for increasing the constraints, second in round-robin mode and the third time without incrementing (by going directly to the full constraint lengths). For each of those runs of the synthesis algorithms, the required computational time and the number of states in the last considered situation graph is presented in the table. Additionally, it is noted how many constraints on $EGO$ were given and if a winning strategy for $EGO$ exists. A question mark “?” in the table stands for data that was not recorded due to hardware limitations (thrashing due to reached memory limits; in fact, this phenomenon could also have lead to the high computational time in experiment number 2 without incrementing).
@@
sequential round-robin
No. Win? $|\mathcal{C}|$ ||states|| time [sec] ||states|| time [sec] ||states|| time [sec]
1 Yes 1 284,636 37 284,636 42 3,692,986 316
@@ 5 Yes 2 537,970 111 152,882 51 ? ?
6 Yes 1 6,297 0.4 6,297 0.4 3,031,519 579
7 Yes 2 1,990,179 709 1,538,446 626 3,351,204 544
The experiments were manually constructed and are not modeling real use cases. Some of the experiments were drawn as grid games that were automatically translated to game graphs. Those grid games were loosely motivated by the challenge of navigating a robot on a factory floor, while the robot should visit predefined areas on a regular basis (defined as counting constraints on $EGO$) and some tiles of the grid might be temporarily blocked by the environment (think e.g. of a fork lifter). To give an impression on the experiments, Figure 10 shows the grid game on which the game graph for Experiment 5 in Table 1 is based on.
mc
Figure 10. Grid game illustrating experiment 5. The circle and the diamond indicate the (initial) positions of $EGO$ and $ALTER$, respectively. The m (machine) and c (charging station) in the upper left and right corners represent target cells of $EGO$: EGOEGO is required to fulfill the counting constraints EGOEGO plays c at least 1 times out of 13 of its own turns.” and EGOEGO plays m at least 1 times out of 13 of its own turns.”, meaning that EGOEGO shall visit the charging station and the machine c at least once in 13 turns, respectively. ALTERALTER has the constraints “”ALTERALTER plays circle at least 1 times out of 1 of its own turns.”” (meaning that $ALTER$ is not allowed to visit the shaded cell in the middle) and ALTERALTER plays (not idle) at least 2 times out of 4 of its own turns.” (meaning: ALTERALTER is only allowed to stay at the same cell instead of moving to another cell in at most 2 out of 4 turns). The game is played as a safety game with crashes of the two players (both players on the same cell) as unsafe states.
The realized comparisons show the great potential of successively enlarging counting constraints, allowing for incomplete graph constructions due to retrieved information on already winnable states of prior graphs instead of encoding the full constraints directly in a graph for strategy synthesis. In all but one of the experiments, the incremental synthesis required significantly less time and memory. Note that the number of states in Table 1 refers to the number of states in the last constructed situation graph. If incrementing is used, additional memory for the states of the already calculated winning regions of previous increments is required. It suffices to remember a minimal set $M$ of states, such that each state of any of the already calculated winning regions is an extension of a state in the minimal set $M$. The number of states in $M$ can be larger than the next situation graph (as can be seen in the example above and this is also the case for experiment 3 in the table). However, the set $M$ is (heuristically) much smaller than the situation graph without incrementing. For example, experiment 3 in Table 1 required to remember 68,937 states of winning regions for sequential increments and 8,401 states for round-robin increments, which is orders of magnitude smaller than the 1,723,345 states in the situation graph without incrementing. As can be seen in Experiment 7, incremental synthesis is not in general better for games with counting constraints, but can be outperformed by classical synthesis without iteration for some games. This can for example happen if the full length of the counting constraints is required to actually find a winning strategy for $EGO$ and the winning region in former increments stays empty. In such cases, the incremental approach only adds a non-beneficial overhead and the only learned additional information is that the full constraint length is actually needed, which does not contribute to runtime savings.
@@ 5. Discussion and Future Work
Towards cooperative games: The presented approach did not consider incrementing over constraints of $ALTER$, but handled them as being fixed. The rationale for this lies in the property ALTERALTER cannot be forced into constraint violations” of games with counting constraints (Definition 2). In general, a game with counting constraints may satisfy the requirement of $ALTER$ always being able to adhere to its counting constraints, only to find the requirement violated for the game with a modified counting constraint as used in the increments. An illustrative example is provided in Figure 11. $ALTER$ has the counting constraint $CC_{min}(ALTER,b,1,3)$, i.e. $ALTER$ plays $b$ at least once in three of its turns. Recall that ALTERALTER cannot be forced into constraint violations” is defined as $ALTER$ is able to enlarge each prefix that satisfies the constraint such that the resulting prefix is also satisfying the constraint. This property is fulfilled when considering the game graph and the constraints $CC_{min}(ALTER,b,1,1)$ or $CC_{min}(ALTER,b,1,3)$. However, it is violated for $CC_{min}(ALTER,b,2,3)$, since the prefix $(1,a,2,\neg b,4,a,5)$ satisfies the constraint [^1] ALTERALTER could play $b$ forever to complete the prefix to an infinite play that satisfies the constraint. This play is not in $G$, but nonetheless is sufficient according to the definition of a prefix satisfying a constraint in Section 2., but there is no possibility for $ALTER$ to still satisfy the constraint with the next turn. Since the definition of a winning strategy relies on the game property of $ALTER$ not being forceable into counting constraint violations, 2 cannot be extended to increments over $ALTER$-constraints. However, such an extension would offer additional potential for more efficient synthesis algorithms.
@@
Symbolic representation: The presented synthesis approach uses an explicit representation of states in the situation graph as arena. However, symbolic synthesis proves to be significantly more efficient than explicit synthesis algorithms for many (but not all) applications [13]. Since the presented approach already has similarities to antichains and the states of the arena have a special structure (representing a snippet of the history of a play), we expect that the approach can be transformed into a symbolic algorithm. We plan to investigate a symbolic version of the algorithm and to compare its performance with its explicit version.
Combining the presented work with on-the-fly synthesis algorithms: Other synthesis algorithms pursuing an on-the-fly forward expansion of a game graph would be a natural point of comparison to the presented approach. In order to apply these algorithms in the setting as in this article, they would however have to be applied to the full game graph with memory for the counting constraints, which they would then expand on-the-fly. We would argue that this is orthogonal to our incremental expansion of the game graph by successively refined counting constraints, and that these on-the-fly algorithms could consequently be combined with our approach. An experimental evaluation however requires substantial implementation effort and is considered as future work. We plan to investigate if the general ideas in the submission can be used for guiding the state space exploration of on-the-fly-methods when facing counting constraints (or other constraints with some monotonicity property) as part of the specification.
Symmetric synthesis: In many application scenarios, the considered system is a system of systems, whose individual systems are identical in their hardware and run in the same context. From the development, production and maintenance point of view, it is preferable to also have the same software and in particular the same controller for all of those systems, as it is typically realized in industrial applications. An example is a fleet of transport robots that take over logistic tasks on a factory floor. Some synthesis approaches that consider such symmetric synthesis (e.g., [7], [3], [21]) exist, but to the best of our knowledge, there still remain open gaps. We are especially interested in such symmetric setting because (1) it is of high relevance for the same application domain where the problem of synthesis with counting constraints arose (fleet of transport robots) and
@@ 6. Conclusion
Synthesis algorithms for reactive systems are promising tools for various engineering tasks, most prominently for the creation of correct-by-construction controllers and for checking the feasibility of specifications. The efficiency of such algorithms is a challenge for getting synthesis into application, since the translation of the system specification into an automaton that is suitable for synthesis is costly in terms of memory and computational time. The exploitation of specific properties in the specification can help to overcome this challenge. In this article, we have shown the potential of incremental synthesis algorithms for specifications with monotonicity properties as for the presented counting constraints. With each increment, the automaton encoding the specification is becoming larger. The key idea is to gather information in each increment that can be used in the next increment to reduce the size of the automaton. The precise nature of this information depends on the considered specification. We have shown an incremental algorithm for a set of counting constraints of the form “the system does a specific move $m$ at least/at most in $k$ turns out of $l$, in which information on winnable states of the automaton in one increment can be used to determine which parts of the automaton for the next increment do not need to be constructed. The approach allows to tackle games with various winning conditions and the integration of specialized synthesis algorithms for calculating the winning regions of intermediate games occurring in the incrementing process. In the majority of the presented experiments, the incremental approach requires significantly less memory and computational time than direct synthesis with full specification translation into one automaton. As future work, we plan to extend the incremental synthesis in five dimensions: (1) Consideration of more cooperative behavior between system and its environment instead of a purely adversarial setting, (2) identification of new specification types with monotonicity properties that can be exploited via incremental synthesis, (3) transformation of the synthesis approach to a symbolic synthesis algorithm, (4) investigation if on-the-fly-methods can benefit from the structural properties inherent in games with counting constraints and (5) consideration of systems of systems, in which identical systems follow the same strategy.